# Sleeping Kelly is a Thirder

### v1

@@ Abstract
The Sleeping Beauty problem was presented by Elga to highlight the role of probabilities in situations with imperfect recall. One approach to solving the Sleeping Beauty problem is to allow Sleeping Beauty to make decisions based on her beliefs, and then characterize what it takes for her decisions to be “rational”. In particular, she can be allowed to make monetary bets based on her beliefs, with the assumption that she wants to gain wealth rather than lose it. However, this approach is often coupled with the assumption that Sleeping Beauty should maximize the expected value of her bets. Here, I argue instead that it is rational for Sleeping Beauty to maximize the growth rate of her wealth using the Kelly Criterion, which leads us to the “thirder” position. Furthermore, this position is shown to be “rational” by Dutch book arguments. If Sleeping Kelly only accepts bets that have a growth rate greater than 1 as a “thirder” then she is not vulnerable to Dutch books. By contrast, if Sleeping Beauty takes the “halfer” position, she is vulnerable to Dutch books.
@@ 1 Problem Statement
2 Background
The two most common answers are to assign probabilities of $\frac{1}{2}$ (“halfer” position) and $\frac{1}{3}$ (“thirder” position) to the coin having landed heads. Elga (2000) argued for the thirder position, while Lewis (2001) prominently argued for the halfer position, but debate continues. According to Bostrom (2007), “neither side seems to have gained a decisive advantage.”
Upon waking Sleeping Beauty, if we ask her to assign a probability to the coin landing heads this leaves open the question of her objective. In other words, how bad is it to say $\frac{1}{3}$ if the “true” answer is $\frac{1}{2}$ and vice versa? Probabilities do not have inherent physical meaning, but objectives can. Allowing Sleeping Beauty to make decisions forces her beliefs to make contact with reality, and we can create criteria to evaluate these decisions. Here enters decision theory.
The simplest type of decision to evaluate is perhaps a monetary bet, in which the basic assumption that Sleeping Beauty would prefer to increase her wealth. The trouble with using monetary bets is that one must make assumptions about Sleeping Beauty’s betting strategy based on the probabilities she assigns to different events. The existing literature from decision theory on the Sleeping Beauty problem uniformly makes the same assumption: that it is rational for Sleeping Beauty to maximize expected value.
@@ 2 Background
Arguably the most substantive analysis of the Sleeping Beauty problem since Elga (2000) is Halpern (2004), which focuses on synchrony and asynchrony. However, these arguments are complicated, and as we will see, as soon as Halpern (2004) considers betting arguments the reasoning falls flat, precisely because it focuses on expectation values of bet payoffs.
In many betting-based arguments, the expected value (ensemble average) of an observable is conflated with the growth rate. If Sleeping Beauty is offered a sequence of gambles, or expects to be offered a sequence of gambles, it is rational for her to maximize the (exponential) growth rate of her wealth, which is not the same as maximizing expected value (Kelly, 1956). But to optimize her growth rate, Sleeping Beauty needs to have some information about what fraction of her wealth she is wagering. [^2] No poker player would ever determine how much to bet on a hand without regard for how much money they have left.
The conflation of expectation value with the growth rate leads to erroneous conclusions, such as that if Sleeping Beauty is offered bets with 1:1 odds on a fair coin flip then, “she will be indifferent between accepting them and not accepting them” (Conitzer, 2015). With any finite amount of initial wealth and any fixed bet size this would cause Sleeping Beauty’s wealth to follow an unbiased random walk over repeated sequences awakenings or repeated experiments, which guarantees eventual bankruptcy (Pearson, 1905).
We now examine a version of the Sleeping Beauty problem with wagers that create multiplicative dynamics rather than additive dynamics. All of Sleeping Beauty’s wagers are expressed as a fraction of the money she has, rather than a fixed bet size. In this situation, Sleeping Beauty maximizes the growth rate of her wealth using the Kelly Criterion (Kelly, 1956).
@@ 3 Problem Setup: Sequential Betting
4 Expected Value versus Growth Rate
@@ 4.1 Expected Value
The idea of betting 100% of your wealth on a coin flip, before being put to sleep or when woken up, should feel wrong. It is. If you run the experiment many times in sequence as a single individual, you will eventually go bankrupt with probability 1. Eventually you will lose a bet, and from then on there is no money left to wager.
4.2 Growth Rate and the Kelly Criterion
Now suppose we are going to repeat the experiment many times in sequence, what is the optimal betting strategy?
@@ 4.2 Growth Rate and the Kelly Criterion
Sleeping Beauty achieves the maximum growth rate for her wealth as a “thirder” who sizes her bets using the Kelly Criterion.
Suppose we run the experiment twice in a row, and the coin lands heads once and tails once. Then your wealth at the end of the two experiments is (1a)(1b)(1+a)(1+b)(1+b)(1-a)(1-b)(1+a)$(1+b)$ times your original wealth.33footnotemark: 3 To maximize this value we must choose $a=0$ and $b=\frac{1}{3}$.
Of course, if we run the experiment only twice we are not guaranteed to get one heads and one tails from the coin flips. However, as the number of experiments tends to infinitely, we expect half of the flips to be heads and half to be tails across the experiments, by the law of large numbers. The wealth multiplier is therefore $\sqrt{(1-a)(1-b)(1+a)(1+b)(1+b)}$ per experiment. Thus, $a=0$, $b=\frac{1}{3}$ is the optimal strategy for long-term wealth maximization. You do not bet any amount before being put to sleep, but each time you are woken it is best to wager $\frac{1}{3}$ of your wealth that the coin landed tails.
Figure 1: Wealth multiplier for wager fractions $a$ and $b$.
This answer at least passes the smell test. One should not bet any fraction of your wealth on a fair coin toss before the experiment begins. Imagine choosing $a>0$ and $b=0$, so you are simply successively betting on successive coin flips. Then by repeating the experiment many times in sequence, your wealth will eventually converge to 0 for any value of $0<a\leq 1$, because (1a)(1+a)<1(1-a)(1+a)<1.
From here on, we will use the wealth multiplier instead of the actual growth rate because it is simpler to work with. We refer to Peters and Adamou (2025) for definitions of wealth multipliers and growth rates.
Inferring Probability from the Kelly Criterion
We can now infer the probability implicitly assigned to the coin landing on heads using the Kelly Criterion, which maximizes the growth rate, or equivalently the multiplier, of wealth under multiplicative dynamics (Thorp, 1975; Kelly, 1956; Peters and Adamou, 2025). Given 1:1 odds on any bet when woken, the Kelly Criterion says to bet a fraction of your wealth b=p(1p)b=p-(1-p), where $p$ is the probability when woken that the coin landed on tails. By setting $b=\frac{1}{3}$, we infer $p=\frac{2}{3}$. In other words, you must implicitly assign a probability $\frac{2}{3}$ to the coin landing tails, and $\frac{1}{3}$ to the coin landing heads, each time you are woken. Similarly, since $a=0$ we infer $p=\frac{1}{2}$ before being put to sleep. Sleeping Kelly is a “thirder.”
@@ 5 Dutch Books
We saw in Section 4.2 that the “thirder” position allows Sleeping Beauty to maximize the growth rate of her wealth when betting according to the Kelly Criterion. In that setup, Sleeping Beauty got to decide what fraction of her wealth to wager, and whether to wager it on heads or tails. We now look at a setting in which Sleeping Beauty’s strategy space is more restricted. Sleeping Beauty will be given an offer to bet on the outcome of the coin flip before she is put to sleep and each time she is woken, but the experimenter will determine the wager, and Sleeping Beauty only has the option to accept or reject the bet. When put back to sleep, she will have no memory of what previous bets she has been offered or whether she accepted them. We require that the same bet be offered whenever she is woken, and that it be determined before the experiment begins, so the offered bet cannot depend on the coin’s outcome or what day it is.
The question we want to answer is whether Sleeping Beauty can be subjected to a Dutch Book. As Conitzer (2015) explains,
“One decision-theoretic approach to answering this question is to attempt to construct a Dutch book to which one of the two positions is vulnerable. A Dutch book is a set of bets that the agent in question would all accept individually, but that together ensure that the agent incurs a strict loss overall. In a diachronic Dutch book, the bets are offered at different times[.]”
@@ 5.1 Dutch Book of Expected Value Maximizer
Suppose that Sleeping Beauty were to accept any bet with a positive expected value. Different Dutch books have been constructed for Sleeping Beauties who are “evidential decision theorists” and “causal decision theorists” that maximize expected value. Consider the following basic example against the halfer position from Hitchcock (2004):
@@
The key flaw here is that a rational sequential betting strategy for Sleeping Beauty cannot be defined without any information about how the size of the bet relates to the amount of money she has available. We want to see whether any Dutch book can be created for a version of Sleeping Beauty who is offered sequential wagers expressed as a fraction of her wealth, and who only accepts bets that increase the growth rate of her wealth, i.e., have a wealth multiplier greater than 1.
5.2 Invulnerability of Sleeping Kelly
We want to know whether Sleeping Beauty can be subjected to a Dutch book if she is only willing to accept bets that individually have a growth rate greater than 1, or equivalently with a wealth multiplier greater than 1; i.e., greater than just sitting on the cash and not betting it.
@@ 5.2 Invulnerability of Sleeping Kelly
Sleeping Kelly cannot be subjected to a Dutch book if she is a “thirder” and only accepts bets with a wealth multiplier greater than 1.
@@
We assume that Sleeping Beauty will not accept a bet unless it has a wealth multiplier greater than one based on the probabilities she assigns. As a “thirder” Sleeping Beauty assigns a probability of $\frac{1}{2}$ to the coin landing tails before being put to sleep and probability $\frac{2}{3}$ to the probability of tails upon being woken.
We now have two conditions for whether Sleeping Beauty will accept a bet offered to her. Before sleeping she requires $(1+\alpha_{OH})^{1/2}(1+\alpha_{OT})^{1/2}>1$. Upon waking she requires $(1+\alpha_{WH})^{1/3}(1+\alpha_{OT})^{2/3}>1$ to accept the bet. We can simplify the exponents to get our two conditions for bet acceptance.
@@
If both of the Dutch book conditions hold, then by multiplying them together, we must also have the condition that the wealth multiplier is less than 1:
@@
However, this is not possible. By reordering the terms, we see
@@
according to our bet acceptance conditions.
Thus, if Sleeping Beauty only accepts bets that individually have a wealth multiplier greater than 1, her overall growth rate must be greater than 1 and she cannot be subjected to a Dutch book. Her wealth can only grow or stay constant (if she rejects the offers). ∎

### v2

# The Sleeping Beauty Problem: Sleeping Kelly is a Thirder

@@ Abstract
The Sleeping Beauty problem is a problem of imperfect recall presented by Elga that has received considerable attention. One approach to solving the Sleeping Beauty problem is to allow Sleeping Beauty to make decisions based on her beliefs, and then characterize what it takes for her decisions to be “rational”. In particular, she can be allowed to make monetary bets based on her beliefs, with the assumption that she wants to gain wealth rather than lose it. However, this approach is often coupled with the assumption that Sleeping Beauty should maximize the expected value of her bets. Here, show that Sleeping Beauty maximizes the expected growth rate of her wealth as a “thirder” sizing bets using the Kelly Criterion under multiplicative dynamics. Furthermore, this position is shown to be impervious to Dutch books. By contrast, the “halfer” position is shown to be vulnerable to Dutch books under similar circumstances.
@@ 1 Problem Statement
2 Halfers, Thirders, and Others
The two most common answers to the Sleeping Beauty problem are to assign probabilities of $\frac{1}{2}$ (“halfer” position) and $\frac{1}{3}$ (“thirder” position) to the coin having landed heads. Elga (2000) argued for the thirder position, while Lewis (2001) prominently argued for the halfer position, but debate continues. According to Bostrom (2007), “neither side seems to have gained a decisive advantage”, and Bostrom argues for a separate answer entirely.
When introducing the problem, Elga (2000) included a basic argument for each position, copied below, before arguing for the thirder position.
“First answer: $\frac{1}{2}$, of course! Initially you were certain that the coin was fair, and so initially your credence in the coin’s landing Heads was $\frac{1}{2}$. Upon being awakened, you receive no new information (you knew all along that you would be awakened). So your credence in the coin’s landing Heads ought to remain $\frac{1}{2}$.”
“Second answer: $\frac{1}{3}$, of course! Imagine the experiment repeated many times. Then in the long run, about $\frac{1}{3}$ of the wakings would be Heads-wakings – wakings that happen on trials in which the coin lands Heads. So on any particular waking, you should have credence $\frac{1}{3}$ that that waking is a Heads-waking, and hence have credence $\frac{1}{3}$ in the coin’s landing Heads on that trial. This consideration remains in force in the present circumstance, in which the experiment is performed just once.”
The halfer argument above has a Bayesian flavor. It presumes that we only update the probabilities we assign to events when our information about those events changes. Naturally, much of the work on the Sleeping Beauty problem revolves around Bayesian arguments. By contrast, the thirder argument above makes no mention of updating beliefs. The thirder argument is derived much more closely to
Some observers have also argued that the problem is ill-posed or indeterminate. The frequentist argument goes roughly as follows. Suppose you were to repeat the experiment many times, flipping a new coin for each experiment, but each time you wake Sleeping Beauty on Monday or Tuesday you simply ask her whether the coin landed Heads or Tails, rather than asking her to assign a probability. If she always says Heads, then her guess will be correct approximately $\frac{1}{3}$ of the times she is asked, but she will get the correct answer in $\frac{1}{2}$ of the experiments. We will argue that the intuition behind this argument is essentially correct, but incomplete.
3 Betting Games
@@
A class of approaches to the Sleeping Beauty problem allow Sleeping Beauty to make decisions with consequences each time she is woken, rather than simply asking her to state probabilities. In general, she can be made to play a game of some kind, and we can assess her performance in the game (Shaw, 2013).
One version of the decision-based approach is to allow Sleeping Beauty to make bets on the coin flip’s outcome before she is put to sleep and each time she is woken (Conitzer, 2015c).
The trouble with betting arguments is that one must make assumptions about Sleeping Beauty’s betting strategy and how it depends on the probabilities she assigns to different events. The existing literature from decision theory on the Sleeping Beauty problem uniformly makes the same assumptions: that it is rational for Sleeping Beauty to either maximize the expected value of each individual bet or maximize the expected value of the sum of her bet payoffs (Arntzenius, 2002; Halpern, 2004; Conitzer, 2015c, a, b). Neither assumption is correct.
The core problem with these assumptions is that they confuse bet payoffs with von Neumann-Morgenstern utilities. The difference is revealed lucidly by the infamous coin flip of  Peters and Gell-Mann (2016)(PGM).
4 The PGM Coin Flip
Suppose you are offered to play a game in which you must wager some amount of money $x$ and then a coin will be flipped. If the coin lands heads, then you will gain 50% for a payoff of $1.5x$, but if the coin lands tails you will lose 40% for a payoff of $0.6x$. Naturally, we assume you would strictly prefer winning the bet and increasing your wealth over not betting, and you would strictly prefer not betting to losing the bet and decreasing your wealth. Is it rational for you to play the game? PGM reveals that the answer is indeterminate.
Suppose we were to say that it is rational for you to maximize the expected value of your payoff. The expected value is $\frac{1.5x+0.6x}{2}=1.05x$, or 1.05 times your initial wager. It is commonly thought that this makes it a worthwhile bet because $1.05>1$. Given this, if you have the opportunity to play the game many, many times, should you? Again, the answer is indeterminate.
Suppose you were to play the PGM coin flip game many times in sequence. You start with an initial wager of $x$. In each round the coin is flipped your wealth grows or shrinks and becomes your wager in the next round, with no opportunity to add or remove money from the game. For example, suppose in the first two rounds you flip heads twice in a row. Then your wealth would be $x(1.5)^{2}=2.25x$. If the coin flips tails twice in a row, then your wealth would be $x(0.6)^{2}=0.36x$. And finally, if you flip heads once and tails once, in either order, then your wealth would be $x(1.5)(0.6)=0.9x$ after two rounds. Notice how after an even number of rounds $2n$, if you flip an equal number of heads and tails, then your wealth has decreased to $x(0.9)^{n}$. You would need to consistently flip more heads than tails just to break even, and you should expect your wealth to decrease at an exponential rate. Playing the sequential game is a bad bet.
Now suppose instead that you were to play the PGM coin flip game many, many times in parallel. After all of the games conclude, you sum your total winnings (or loses) to get your final payoff. Let $2n$ be the number of games you play. In the limit as $n\rightarrow\infty$, you expect the coin to land heads in approximately half of the games and tails in approximately half of the games. Your expected total payoff therefore converges to $x\frac{1.5n+0.6n}{2n}=1.05x$. In other words, the expected value of each individual stage game is the ensemble average over many parallel plays of the game in the infinite limit. In this scenario, playing the ensemble of games is a great bet, at least for sufficiently large $n$.
Ultimately, when asked about a single stage game – one bet on one coin flip – you cannot determine whether or not the bet is worth taking without making assumptions about the broader dynamics of your wealth – what bets you are taking in parallel, if any, and what bets you may make in the future, if any.
One of the key features of the Sleeping Beauty problem, and other problems of imperfect recall, is that events occur in sequence. To use betting arguments for these problems, we need to be careful in specifying the appropriate wealth dynamics. This is precisely where existing Dutch Book arguments for the Sleeping Beauty problem go wrong. We now examine the Sleeping Beauty problem in which she is allowed to make bets each time she is awakened, recognizing that these actions take place in sequence.
@@ 5 Problem Setup: Sequential Betting
6 Expected Value versus Wealth Multiplier
@@ 6.1 Expected Value
The idea of betting 100% of your wealth on a coin flip, before being put to sleep or when woken up, should feel wrong. It is. If you run this experiment many times in sequence, where your wealth at the end of each experiment becomes your wealth at the start of the next, then this strategy guarantees you eventually go bankrupt (with probability 1). Eventually you will lose a bet, and from then on there is no money left to wager.
Similarly, if we consider using this strategy in the “Extreme Sleeping Beauty” version of the experiment from Bostrom (2007) where Sleeping Beauty is awakened 1 million days in a row when the coin lands tails rather than only twice, bankruptcy is almost assured in even a single experiment.
6.2 Wealth Multiplier
If we do plan to repeat the experiment many times in sequence, where the final wealth after one experiment becomes the initial wealth at the next, what is Sleeping Beauty’s optimal betting strategy?
Imagine we run the experiment twice in a row, and the coin lands heads once and tails once. Then your wealth at the end of the two experiments is (1a)(1b)(1+a)(1+b)(1+b)(1-a)(1-b)(1+a)$(1+b)$ times your original wealth.33footnotemark: 3 To maximize this value we must choose $a=0$ and $b=\frac{1}{3}$.
Of course, if we run the experiment only twice we are not guaranteed to get one heads and one tails from the coin flips. However, as the number of experiments tends to infinitely, we expect half of the flips to be heads and half to be tails across the experiments, by the law of large numbers. The wealth multiplier is therefore $\sqrt{(1-a)(1-b)(1+a)(1+b)(1+b)}$ per experiment, and ($a=0$, $b=\frac{1}{3}$) is the optimal strategy for long-term wealth maximization. You do not bet any amount before being put to sleep, but each time you are woken it is best to wager $\frac{1}{3}$ of your wealth that the coin landed tails.
Figure 1: Wealth multiplier for wager fractions $a$ and $b$.
This betting strategy at least passes the smell test. One should not bet any fraction of your wealth on a fair coin toss before the experiment begins. Imagine choosing $a>0$ and $b=0$, so you are simply betting on successive coin flips. Then by repeating the experiment many times in sequence, your wealth will converge to 0 in the limit for any value of $0<a\leq 1$, because (1a)(1+a)<1(1-a)(1+a)<1.
We refer to Peters and Adamou (2025) for formal definitions of wealth multipliers and time-average growth rates. And we emphasize that the wealth multiplier, which is the value whose expectation we want to maximize, is unit-less unlike the payoff expectation value from earlier which was in units of dollars.
Now, given the strategy ($a=0$, $b=\frac{1}{3}$) for maximizing Sleeping Beauty’s wealth multiplier, we want to infer the probability she implicitly assigns to the coin having landed heads each time she is woken.
Inferring Probability from the Kelly Criterion
Recall that in our sequential betting problem, before she is put to sleep and each time she is woken, Sleeping Beauty gets to wager any fraction of her wealth at 1:1 odds. So for each single bet, the amount that she stands to win is the same amount she stands to lose. The Kelly criterion is the betting strategy that maximizes the long-term expected geometric growth rate of one’s wealth, which in our case corresponds to maximizing the wealth multiplier for a sequence of bets. With 1:1 odds, the Kelly criterion says to wager a wealth fraction b=p(1p)b=p-(1-p), where $p$ is the probability when woken that the coin landed on tails (Thorp, 1975; Kelly, 1956).
By setting $b=\frac{1}{3}$, we infer $p=\frac{2}{3}$. In other words, you must implicitly assign a probability $\frac{2}{3}$ to the coin landing tails, and $\frac{1}{3}$ to the coin landing heads, each time you are woken. Similarly, since $a=0$ we infer $p=\frac{1}{2}$ before being put to sleep. Sleeping Kelly is a thirder.
@@ 6.2 Wealth Multiplier
Sleeping Beauty maximizes the growth rate of her wealth as a thirder who sizes bets according to the Kelly criterion.
We have invoked the Kelly criterion here because it translates probabilities into bet sizes that maximize the wealth multiplier, and we are using wealth multiplier maximization as our choice criterion. However, this is not the only choice criterion Sleeping Beauty should care about. Sleeping Beauty must also avoid making herself vulnerable to Dutch Books.
@@ 7 Dutch Books
So far, we have seen that the thirder position allows Sleeping Beauty to maximize the growth rate of her wealth when sizing bets according to the Kelly Criterion. In that setup, Sleeping Beauty got to decide what fraction of her wealth to wager, and whether to wager it on heads or tails. We now look at a setting in which Sleeping Beauty’s strategy space is more restricted. Sleeping Beauty will be given an offer to bet on the outcome of the coin flip before she is put to sleep and each time she is woken, but the experimenter will determine the wager, and Sleeping Beauty only has the option to accept or reject the bet. When put back to sleep, she will have no memory of what previous bets she has been offered or whether she accepted them. We require that the same bet be offered whenever she is woken, and that it be determined before the experiment begins, so the offered bet cannot depend on the coin’s outcome or what day it is or Sleeping Beauty’s prior decisions. In other words, the experimenter can never make choices about what bets to offer using information that is not available to Sleeping Beauty.
The question we want to answer is whether Sleeping Beauty can be subjected to a diachronic Dutch Book. As Conitzer (2015b) explains,
“A Dutch book is a set of bets that the agent in question would all accept individually, but that together ensure that the agent incurs a strict loss overall. In a diachronic Dutch book, the bets are offered at different times[.]”
@@ 7.1 Dutch Book of Expected Value Maximizer
Different Dutch books have been constructed for Sleeping Beauties who are “evidential decision theorists” and “causal decision theorists” that maximize the expected value of their bets (Conitzer, 2015a, b). Consider the following basic example against the halfer position from Hitchcock (2004):
@@
The key flaw here is assuming that a rational Sleeping Beauty would accept any of the individual bets in this example. Observe that if you start with any finite amount of wealth and place a sequence of wagers of the same size one a fair coin with 1:1 payoff odds, then your wealth dynamics follow an unbiased random walk which guarantees eventual bankruptcy. Perturbing the payoff odds by some small $\epsilon$ to give a slightly biased random walk still leaves an overwhelming chance of ruin.
A rational sequential betting strategy for Sleeping Beauty cannot be defined independent of the fraction of her wealth being put at stake. The relation between the size of her bet and current endowment of wealth cannot be ignored when considering her wealth dynamics. We want to see whether any Dutch book can be created for a Sleeping Beauty who only accepts bets with a wealth multiplier greater than 1.
7.2 Invulnerability of Sleeping Kelly as a Thirder
We want to know whether Sleeping Beauty can be subjected to a Dutch book if she is only willing to accept bets that individually have a wealth multiplier greater than 1; i.e., greater than just sitting on the cash and not betting it.
@@ 7.2 Invulnerability of Sleeping Kelly as a Thirder
Sleeping Kelly cannot be subjected to a Dutch book if she is a thirder and only accepts bets with a wealth multiplier greater than 1.
@@
Recall that as a thirder Sleeping Beauty assigns a probability of $\frac{1}{2}$ to the coin landing tails before being put to sleep and probability $\frac{2}{3}$ to the probability of tails upon being woken. We now have two conditions for whether Sleeping Beauty accepts a bet. Before sleeping she requires $(1+\alpha_{OH})^{1/2}(1+\alpha_{OT})^{1/2}>1$ and upon waking she requires $(1+\alpha_{WH})^{1/3}(1+\alpha_{OT})^{2/3}>1$. We can simplify the exponents to get our two conditions for bet acceptance.
@@
If both of the Dutch book conditions hold, then by multiplying them together, we see that the wealth multiplier must be less than 1:
@@
However, this is not possible according to our bet acceptance conditions. By reordering the terms, we see
@@
Thus, Sleeping Kelly cannot be subjected to Dutch books if she is a thirder. ∎